OpenStudy (osuxbucks15):

Thermochemical equation..... Calculate the calories required to convert 84.2g of H2O(s) at 0.0 degrees C to H2O(g) at 100 degrees C.

OpenStudy (mertsj):

How many degrees do we want to raise the water?

OpenStudy (osuxbucks15):

100?

OpenStudy (mertsj):

how many grams do we want to heat?

OpenStudy (osuxbucks15):

84.2g

OpenStudy (mertsj):

How many calories does it take to raise 1 g of water 1 degree?

OpenStudy (osuxbucks15):

is it 4?

OpenStudy (mertsj):

What is the definition of a calorie?

OpenStudy (osuxbucks15):

ummm, Im not sure, i no it equals 4.18 J if we need that for this?

OpenStudy (osuxbucks15):

Either of two units of heat energy.

OpenStudy (osuxbucks15):

84.2g x 100 x 4/ 18?

OpenStudy (anonymous):

there are two things involved here 1) raise temperature of water from 0 to 100 degrees -> use the specific heat for water 2) convert liquid water at 100deg to gas at 100deg -> use latent heat of vaporization

OpenStudy (anonymous):

total heat = sum of heat for both the steps

OpenStudy (anonymous):

wait.. another one step 0) convert SOLID water at 0 to liquid water at 0 -> latent heat of fusion

OpenStudy (anonymous):

total heat = sum of all three heats

OpenStudy (osuxbucks15):

336.8/ 84.2g x 100 for step 1 = .04

OpenStudy (anonymous):

start with step 0

OpenStudy (osuxbucks15):

ok

OpenStudy (anonymous):

us the equation button below so I can understand.

OpenStudy (osuxbucks15):

o ok never saw that before, and im not sure how to do step 0

OpenStudy (anonymous):

same exact thing like step 2 just a different value for Lf

OpenStudy (osuxbucks15):

can u give me the equation, and ill plug in the numbers, because i dont know the equation

OpenStudy (anonymous):

\[Q_0=mL_f=84.2{\rm g}\times L_f\] Lf is the latent heat of fusion.. should be in a table in your book

OpenStudy (osuxbucks15):

506.0?

OpenStudy (osuxbucks15):

84.2 x 6.01

OpenStudy (anonymous):

you cannot solve without the unitz

OpenStudy (anonymous):

84.2 g 6.01 what?

OpenStudy (osuxbucks15):

84.2g x 6.01 H2O

OpenStudy (osuxbucks15):

\[H _{2}O\]

OpenStudy (anonymous):

H2O is the chemical formula for water. It is not a unit for a physical quantity units are like.. grams for mass, second for time, calori or joule for heat, etc..

OpenStudy (anonymous):

you are looking for the latent heat of fusion.

OpenStudy (anonymous):

i.e., the amount of heat required to convert 1g of ice to water.

OpenStudy (osuxbucks15):

delta H\[_{fus}\] (KJ/mol) = H2O 6.01

OpenStudy (anonymous):

that is the latent heat of fution for ice or water.

OpenStudy (anonymous):

ok. so, this value is in moles and the amount we have is in grams so, first convert 84.2g to moles for water

OpenStudy (anonymous):

how mny moles of water do we have in 84.2g?

OpenStudy (osuxbucks15):

84.2g x 6.01 x \[10^{23}\]

OpenStudy (osuxbucks15):

= 5.06 x \[10^{25}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1mole of water = 2(1)+16=18g

OpenStudy (osuxbucks15):

4.6 moles?

OpenStudy (anonymous):

so, how many moles do we have in 84.2g?

OpenStudy (anonymous):

good lets use 4.68

OpenStudy (osuxbucks15):

84.2/18= 4.7

OpenStudy (osuxbucks15):

o ok 4.68 works to

OpenStudy (anonymous):

now ou can multiply apples with apples.

OpenStudy (osuxbucks15):

28.1

OpenStudy (anonymous):

\[Q_0=4.68{\rm mol}\times6.01{\rm kJ/mol}\] =?

OpenStudy (anonymous):

28.1 what ? joules, kilo joules, bananas, calories,...?

OpenStudy (osuxbucks15):

28.1 KJ/mol

OpenStudy (anonymous):

we have to be careful or evrything will go waste

OpenStudy (anonymous):

notice that the mol gets cancelled. so, \[Q_0=28.1kJ\]

OpenStudy (anonymous):

so, we apply this much heat and all that ice will turn to water

OpenStudy (anonymous):

now we heat the water up to 100deg \[Q_1=mc\Delta T\]

OpenStudy (osuxbucks15):

not sure how to do that... 28.1KJ x 100 deg?

OpenStudy (anonymous):

m=84.2g c=4.186J/g delta T = 100-0=100

OpenStudy (osuxbucks15):

35246.1 J/ Degrees C

OpenStudy (osuxbucks15):

g degrees C

OpenStudy (anonymous):

heat is just joules or KJ or cal or Kcal no more things with it. so, \[Q_1=35246J=35.25kJ\] since our first one was in KJ, we make all our heats to KJ

OpenStudy (anonymous):

now, the liquid water is ready to be turned to steam and blow up:)

OpenStudy (anonymous):

so, we do Latent Hat for Vaporization. \[Q_3=mL_v\]

OpenStudy (osuxbucks15):

4.68 x 40.7 Delta H vap ( KJ/mol)

OpenStudy (osuxbucks15):

4.68 moles

OpenStudy (osuxbucks15):

190.5KJ

OpenStudy (anonymous):

nice work

OpenStudy (anonymous):

I cannot verify the values for the latent heats but thats the steps to do this problem.

OpenStudy (anonymous):

so, total heat for the reaction = sum of all the heats \[Q=Q_0+Q_1+Q_2\]

OpenStudy (anonymous):

when you add, make sure they all have the same units. (in our case, it is kJ)

OpenStudy (anonymous):

the final heat will also be in kJ

OpenStudy (osuxbucks15):

what is latent heat?

OpenStudy (anonymous):

but I think the question asks for the value in calories... so do the necesary conversion

OpenStudy (anonymous):

Latent heat is the amount of heat required to change the physical state of a unit mass of a subtance at a constant temperature.

OpenStudy (osuxbucks15):

so where would i find latent heat, in another table?

OpenStudy (anonymous):

like from solid to liquid at its melting point or from liquid to solid at melting point from liquid to gas at boiling point and backwards

OpenStudy (osuxbucks15):

what elements or things do i need to find them for?

OpenStudy (anonymous):

any element or compound that can change the states will have the corresponding latent heat

OpenStudy (osuxbucks15):

ok, im sorry but im really confused right now

OpenStudy (anonymous):

what with?

OpenStudy (osuxbucks15):

i have no idea how to change latent heat or get it

OpenStudy (anonymous):

it will be provided to you in the test. In homeworks, they expect you to use the textbook (as they usually have the values for common materials)

OpenStudy (anonymous):

you do not change latent heats. they are the properties of that material. you have to take it as given

OpenStudy (osuxbucks15):

4.18 for water?

OpenStudy (anonymous):

4.18J/gC is the "specific heat for water"

OpenStudy (osuxbucks15):

Latent heat of melting - 334 kJ/kg

OpenStudy (osuxbucks15):

for water

OpenStudy (osuxbucks15):

its not negative the dash is like an equal sign

OpenStudy (anonymous):

right.. units are different but representing the same quantity.

OpenStudy (osuxbucks15):

now what?

OpenStudy (anonymous):

did you add up the three heats??

OpenStudy (osuxbucks15):

what other heats do i need to find? for what elements would it be 190.5 and 32.25?

OpenStudy (anonymous):

??? are we still talking about this question or jumping out randomly?

OpenStudy (osuxbucks15):

no thats what we got before, im not sure what i need to find the heat for now?

OpenStudy (anonymous):

ADD em up

OpenStudy (anonymous):

the three heats that we found

OpenStudy (osuxbucks15):

wait, im just wondering are those the heats?

OpenStudy (anonymous):

yes the "Q" is the symbol for heats

OpenStudy (anonymous):

kJ is the unit of measurement for heat

OpenStudy (osuxbucks15):

o ok

OpenStudy (osuxbucks15):

so i got 32.25 + 190.5 + 334= 556.8KJ